# Generalized Out-of-Distribution Detection: A Survey

### v1

@@ 2.4 Out-of-Distribution Detection
Background  With the observation that deep learning models can overconfidently classify samples from different semantic distributions, the field of out-of-distribution detection emerges, requiring models to reject label-shifted samples to guarantee reliability and safety.
Definition  Out-of-distribution detection aims to detect test samples with non-overlapping labels w.r.t training data. Formally, test samples in the OOD detection setting come from the distribution with semantic shift from ID, i.e., $P(Y)\neq P^{\prime}(Y)$. The ID can contain a single class or multiple classes. When multiple classes exist in training, OOD detection should NOT harm the ID classification capability.
Position in Framework  Out-of-distribution detection is a super-category that includes semantic AD, one-class ND, multi-class ND, and open-set recognition. In the multi-class setting, the problem can be canonical to OSR (Section 4)—accurately classify test samples from ID within the class space $\mathcal{Y}$, and reject OOD test samples with semantics outside the support of $\mathcal{Y}$. However, OOD detection encompasses a broader spectrum of learning tasks (e.g., multi-label classification [55], reinforcement learning) and solution space (e.g., density estimation and outlier exposure). Some approaches relax the constraints imposed by OSR and achieve strong performance. Moreover, OOD detection also includes other sub-topics of one-class novelty detection and semantic anomaly detection where a single class exists in the in-distribution.
Application and Benchmark  The application of OOD detection usually falls into safety-critical situations such as autonomous driving [56, 57]. An example academic benchmark is to use CIFAR-10 as ID during training, and to distinguish CIFAR images from other datasets such as SVHN, etc. Researchers should pay attention that OOD datasets should NOT have label overlapping with ID datasets when building the benchmark.
@@
Remark  While most works in the current community interpret the keyword ”out-of-distribution” as ”out-of-label/semantic-distribution”, some OOD detection works also consider detecting covariate shifts [58], which claim that covariate shift usually leads to a significant drop in model performance and therefore needs to be identified and rejected. However, although detecting covariate shift is reasonable on some specific (usually high-risk) tasks, such as a medical diagnosis model that trained by one hospital should detect scans under distributional shift, research on this topic remains a controversial task w.r.t OOD generalization tasks (c.f. Section 2.6). Detecting semantic shift has been the mainstream of OOD detection tasks.
@@ 2.5 Outlier Detection
2.6 Discussion
In this subsection, we further contrast and summarize how the five sub-topics described above fit in our generalized OOD detection framework. As shown in Figure 1, semantic AD and one-class ND have identical problem formulation, despite the subtle difference in motivation. Multi-class ND and OSR both focus on semantic shift w.r.t a multi-class classification model. The only difference is that multi-class ND does not require ID classification whereas OSR does.
Despite the practical relevance of OSR, several restrictions remain such as no additional data is permitted during training and a required guarantee on theoretical open-risk bound. These restrictions exclude methods that focus more on effective performance improvement but may violate the constraints of OSR. On the other hand, OOD detection encompasses a broader spectrum of learning tasks and solution space (to be discussed in Section 5). Some approaches relax the constraints imposed by OSR and achieve strong performance.
@@ 2.6 Discussion
Interestingly, the outlier detection task can be considered as an outlier in the generalized OOD detection framework, since outlier detectors are operated on the scenario when all observations are given, rather than following the training-test scheme. Also, publications exactly on this topic are rarely seen in the recent deep learning venues. However, we still include outlier detection in our framework, because intuitively speaking, outliers also belong to one type of out-of-distribution, and introducing it can help familiarize readers more with various terms (e.g., OD, AD, ND, OOD) that have confused the community for a long while. Additionally, we briefly discuss five related topics below, and further clarify the scope of this survey.
@@
Learning with Rejection can date back to early works on abstention [71, 72], which considered simple model families such as SVMs [73]. The phenomenon of neural networks’ overconfidence in out-of-distribution data is first revealed by [74]. Despite methodologies differences, subsequent works developed on OOD detection and OSR share the underlying spirit of classification with rejection option.
@@
Domain Adaptation (DA) [12] and Domain Generalization (DG) [75] also follow “open-world” assumption. Different from generalized OOD detection settings, DA/DG expects the existence of covariate shift during testing without any semantic shift, and requires classifiers to make accurate predictions regardless of the covariate shift [76]. Noticing that OOD detection commonly concerns detecting the semantic shift, which is complementary to DA/DG. In the case when both covariate and semantic shift take place, the model should be able to detect semantic shift while being robust to covariate shift. More discussion on relations between DA/DG and OOD detection is in Section 7.2. The difference between DA and DG is that while the former requires extra but few training samples from the target domain, the latter one does not.
@@
Novelty Discovery [77] requires all observations are given in advance as outlier detection does. The observations are provided in a semi-supervised manner, and the goal is to explore and discover the new categories and classes in the unlabeled set. Different from outlier detection where outliers are sparse, the unlabeled set in novelty discovery setting can mostly consist of, and even be overwhelmed by unknown classes.
@@
Zero-shot Learning [78] has a similar goal of novelty discovery, but follows the training-testing scheme. The test set is under the “open-world” assumption with unknown classes, which expect classifiers trained only on the known classes to perform classification on unknown testing samples with the help of extra information such as label relationships.
@@
Open-world Recognition [79] aims to build a lifelong learning machine that can actively detect novel images [80], label them as new classes, and perform continuous learning. It can be viewed as the combination of novelty detection and incremental learning.
@@
Anomaly Detection
One-Class
Novelty Detection
@@
Density
$\S$~3.1.1: Classic Density Est. [81, 82, 83, 84, 85, 86]
$\S$~3.1.2: NN-based Density Est. [87, 88, 89, 90, 91, 92, 93, 89, 94, 95, 96, 97, 98]
$\S$~3.1.3: Energy-based Models [99, 100, 101]
$\S$~3.1.4: Frequency-based Methods [102, 103, 104, 105]
$\S$~3.2
Reconstruction
$\S$~3.2.1: Sparse Representation [106, 107, 108, 109, 110]
$\S$~3.2.2: Reconstruction-Error
[111, 112, 113, 114, 111, 112, 115, 116, 117, 118, 89, 119, 120, 113, 121, 122, 123, 124]
$\S$~3.3
@@
$\S$~3.3.1: One-Class Classification [125, 126, 127, 128]
$\S$~3.3.2: PU Learning
@@
$\S$~3.3.3: Self-Supervised Learning [142, 143, 144, 145, 146]
$\S$~3.4: Distance-based Methods [147, 148, 149]
$\S$~3.5: Gradient-based Methods [150]
$\S$~3.6: Discussion and Theoretical Analysis [151, 152]
@@
Out-of-Distribution
Detection
@@
$\S$~5.1.1.a: Post-hoc Detection [187, 188, 189, 55, 190, 191, 192]
$\S$~5.1.1.b: Conf. Enhancement
[193, 194, 195, 196, 197, 198, 199, 200, 201, 202, 203, 204, 205, 206, 207, 192, 208, 209, 58]
$\S$~5.1.1.c: Outlier Exposure (OE) [210, 52, 211, 212, 213, 214, 215, 216, 217, 218, 219]
$\S$~5.1.2: OOD Data Generation [220, 221, 222, 223]
$\S$~5.1.3: Gradient-based Methods [188, 191]
$\S$~5.1.4: Bayesian Models [224, 225, 226, 227, 228, 229]
$\S$~5.1.5: Large-scale OOD Detection [171, 230, 231, 168]
$\S$~5.2: Density-based Methods
[87, 88, 89, 90, 121, 207, 92, 232, 233, 234, 235, 236, 237, 238, 239, 240]
$\S$~5.3: Distance-based Methods [207, 241, 242, 243, 244, 245, 246]
@@
3 Anomaly Detection & One-Class Novelty Detection: Methodology

### v2

@@ 2.4 Out-of-Distribution Detection
Background  With the observation that deep learning models are often inappropriate but in fact overconfident in classifying samples from different semantic distributions in the image classification task and text categorization [64], the field of out-of-distribution detection emerges, requiring the model to reject inputs that are semantically different from the training distribution and therefore should not be predicted by the model.
Definition  Out-of-distribution detection, or OOD detection, aims to detect test samples that drawn from a distribution that is different from the training distribution, with the definition of distribution to be well-defined according to the application in the target. For most machine learning tasks, the distribution should refer to “label distribution”, which means that OOD samples should not have overlapping labels w.r.t. training data. Formally, in the OOD detection, the test samples come from a distribution whose semantics are shifted from ID, i.e., $P(Y)\neq P^{\prime}(Y)$. Note that the training set usually contains multiple classes, and OOD detection should NOT harm the ID classification capability.
Position in Framework  Out-of-distribution detection can be canonical to OSR in common machine learning tasks like multi-class classification—keeping the classification performance on test samples from ID class space $\mathcal{Y}$, and reject OOD test samples with semantics outside the support of $\mathcal{Y}$. Also, the multi-class setting and the requirement of ID classification distinguish the task from AD and ND.
Application and Benchmark  The application of OOD detection usually falls into safety-critical situations, such as autonomous driving [65, 66]. An example academic benchmark is to use CIFAR-10 as ID during training, and to distinguish CIFAR images from other datasets such as SVHN, etc. Researchers should pay attention that OOD datasets should NOT have label overlapping with ID datasets when building the benchmark.
@@
Remark: OSR vs. OOD Detection  The difference between OSR and OOD detection tasks are three-fold.
1) Different benchmark setup: OSR benchmarks usually split one multi-class classification dataset into ID and OOD parts according to classes, while OOD detection takes one dataset as ID and find several other datasets as OOD with the guarantee of non-overlapping categories between ID / OOD datasets. However, despite the different benchmark traditions of the two sub-tasks, they are in fact tackling the same problem of semantic shift detection.
2) No additional data in OSR: Due to the requirement of theoretical open-risk bound guarantee, OSR discourages the usage of additional data during training by design [28]. This restriction precludes methods that are more focused on effective performance improvements (e.g., outlier exposures [67]) but may violate OSR constraints.
3) Broadness of OOD detection: Compare to OSR, OOD detection encompasses a broader spectrum of learning tasks (e.g., multi-label classification [68]), wider solution space (to be discussed in Section 3).
Remark: Mainstream OOD Detection Focuses on Semantics  While most works in the current community interpret the keyword ”out-of-distribution” as ”out-of-label/semantic-distribution”, some OOD detection works also consider detecting covariate shifts [69], which claim that covariate shift usually leads to a significant drop in model performance and therefore needs to be identified and rejected. However, although detecting covariate shift is reasonable on some specific tasks (usually due to high-risk or privacy reasons) that to be discussed in the following paragraph, research on this topic remains a controversial task w.r.t OOD generalization tasks (c.f. Section 2.6 and Section 7.2). Detecting semantic shift has been the mainstream of OOD detection tasks.
Remark: Extension of OOD Detection  We provide another definition from the perspective of generalization: Out-of-distribution detection, or OOD detection, aims to detect test samples to which the model cannot or does not want to generalize [70]. In the most of the machine learning tasks, such as image classification, the models are expected to generalize their prediction capability to samples with covariate shift, and they are only unable to generalize when semantic shift occurs. However, for applications where models are by-design nontransferable to other domain, such as many deep reinforcement learning tasks like game AI [71, 72], the key term “distribution” should refer to “data / input distribution”, so that the model should refuse to make decision under the environment that is not the same as training environment, i.e., $P(X)\neq P^{\prime}(X)$. Similar applications are those high-risk tasks such as medical image classification [73] or in privacy-sensitive scenario [74], where the models are expected to be very conservative and only make predictions for samples exactly from the training distribution, rejecting any samples that deviate from it. In sum, an OOD detection task can be valid when the “detection” task has well reconciled with the “generalization” task. By all means, detecting semantic shift is still the mainstream of OOD detection tasks and is the focus of this survey.
@@ 2.5 Outlier Detection
Remark: On Inclusion of Outlier Detection  Interestingly, the outlier detection task can be considered as an outlier in the generalized OOD detection framework, since outlier detectors are operated on the scenario when all observations are given, rather than following the training-test scheme. Also, publications exactly on this topic are rarely seen in the recent deep learning venues. However, we still include outlier detection in our framework, because intuitively speaking, outliers also belong to one type of out-of-distribution, and introducing it can help familiarize readers more with various terms (e.g., OD, AD, ND, OOD) that have confused the community for a long while.
2.6 Related Topics
Apart from the five sub-topics that are described in our generalized OOD detection framework (shown in Figure 1), we further briefly discuss five related topics below, which helps clarify the scope of this survey.
@@ 2.6 Related Topics
Learning with Rejection (LWR)  LWR can date back to early works on abstention [87, 88], which considered simple model families such as SVMs [89]. The phenomenon of neural networks’ overconfidence in OOD data is first revealed by [90]. Despite methodologies differences, subsequent works developed on OOD detection and OSR share the underlying spirit of classification with rejection option.
@@
Domain Adaptation / Generalization  Domain Adaptation (DA) [15] and Domain Generalization (DG) [91] also follow “open-world” assumption. Different from generalized OOD detection settings, DA/DG expects the existence of covariate shift during testing without any semantic shift, and requires classifiers to make accurate predictions into the same set of classes [92]. Noticing that OOD detection commonly concerns detecting the semantic shift, which is complementary to DA/DG. In the case when both covariate and semantic shift take place, the model should be able to detect semantic shift while being robust to covariate shift. More discussion on relations between DA/DG and OOD detection is in Section 7.2. The difference between DA and DG is that while the former requires extra but few training samples from the target domain, the latter one does not.
@@
Novelty Discovery  Novelty discovery [93, 94, 95, 96, 97] requires all observations are given in advance as outlier detection does. The observations are provided in a semi-supervised manner, and the goal is to explore and discover the new categories and classes in the unlabeled set. Different from outlier detection where outliers are sparse, the unlabeled set in novelty discovery setting can mostly consist of, and even be overwhelmed by unknown classes.
@@
Zero-shot Learning  Zero-shot learning [98] has a similar goal of novelty discovery, but follows the training-testing scheme. The test set is under the “open-world” assumption with unknown classes, which expect classifiers trained only on the known classes to perform classification on unknown testing samples with the help of extra information such as label relationships.
@@
Open-world Recognition  Open-world recognition [99] aims to build a lifelong learning machine that can actively detect novel images [100], label them as new classes, and perform continuous learning. It can be viewed as the combination of novelty detection and incremental learning.
2.7 Organization of Remaining Sections
In the following sections, we will introduce the methodologies of AD, ND, OSR, OOD detection, and OD, each sub-task in one section. For each sub-task, we categorize and introduce the methodologies into the following four groups: 1) classification-based methods: methods that largely rely on classifiers; 2) density-based methods: detecting OOD by modeling data density; 3) distance-based methods: using distance metrics (usually in the feature space) to identify OODs; and 4) reconstruction-based methods: methods featured by reconstruction techniques. Our following review mainly focuses on summarizing the main ideas of the methodological categories, hoping to provide readers with big pictures for solving OOD-related problems and inspire readers to develop more effective methods. A concurrent survey [101] on reviewing methods under the generalized OOD detection framework provides a detailed explanation of OOD-related methods, greatly complementing our work.
Remark: This Survey Highlights OOD Detection Methods  Notice the existence of comprehensive surveys on AD, ND, and OD methodologies in recent years [20, 21, 22, 23], this survey will highlight the overview of OSR and OOD detection methods in complementing the good surveys mentioned above, so that the ordering of the following methodological sections are: OOD detection in Section 3, OSR in Section 4, AD / ND in Section 5, and OD in Section 6.
@@ 2.7 Organization of Remaining Sections
Out-of-Distribution
Detection
@@
$\S$~3.1.1
Output
-based
Methods
a: Post-hoc Detection [64, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111]
b: Conf. Enhancement
[112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 69, 125]
c: Outlier Exposure
@@ Methods
$\S$~3.1.2: Label Space Redesign [137, 138, 139, 140, 141]
$\S$~3.1.3: OOD Data Generation [142, 143, 144, 145, 146, 147]
$\S$~3.1.4: Gradient-based Methods [102, 148]
$\S$~3.1.5: Bayesian Models [149, 150, 151, 152, 153, 154]
$\S$~3.1.6: Large-Scale OOD Detection [138, 140, 155, 156]
$\S$~3.2: Density-based Methods
[157, 158, 159, 160, 161, 103, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172]
$\S$~3.3: Distance-based Methods [103, 173, 174, 175, 176, 177, 178, 110, 179]
$\S$~3.4: Reconstruction-based Methods [180, 181, 182]
@@
Anomaly Detection
Novelty Detection
@@
Density
$\S$~5.1.1: Classic Density Est. [211, 212, 213, 214, 215, 216]
$\S$~5.1.2: NN-based Density Est.
[157, 158, 159, 160, 217, 162, 218, 159, 219, 220, 221, 222, 223]
$\S$~5.1.3: Energy-based Models [224, 225, 226]
$\S$~5.1.4: Frequency-based Methods [227, 228, 229, 230]
$\S$~5.2
Reconstruction
$\S$~5.2.1: Sparse Representation [231, 232, 233, 234, 235]
$\S$~5.2.2: Reconstruction-Error
[236, 237, 238, 239, 236, 237, 240, 241, 242, 243, 159, 244, 245, 238, 161, 56, 246, 247]
$\S$~5.2.3: Gradient Representation [248]
$\S$~5.3: Distance-based Methods [249, 250, 251]
$\S$~5.4
@@
$\S$~5.4.1: One-Class Classification [252, 253, 254, 255]
$\S$~5.4.2: PU Learning
[256, 257, 258, 259, 260, 261, 262, 263, 264, 265, 266, 267, 268]
$\S$~5.4.3: Self-Supervised Learning [269, 270, 271, 272, 273]
$\S$~5.5: Discussion and Theoretical Analysis [60, 61]
@@
Figure 3: Timeline for representative methodologies of (a) OOD detection, details in Section 3, (b) open set recognition, details in Section 4, and (c) anomaly detection & novelty detection, details in Section 5, under generalized OOD detection framework. Different colors indicate different categories of methodologies. Each method has its corresponding reference (inconspicuous white) in the lower right corner. Methods with open-source code are prioritized for inclusion in this figure.
3 OOD Detection: Methodology
3.1 Classification-based Methods
3.1.1 Output-based Methods
a. Post-hoc Detection  Post-hoc methods have the advantage of being easy to use without modifying the training procedure and objective. The property can be important for the adoption of OOD detection methods in real-world production environments, where the overhead cost of retraining can be prohibitive. Early work ODIN [102] is a post-hoc method that uses temperature scaling and input perturbation to amplify the ID/OOD separability. Key to the method, a sufficiently large temperature has a strong smoothing effect that transforms the softmax score back to the logit space—which effectively distinguishes ID vs. OOD. Note that this is different from confidence calibration, where a much milder $T$ is employed. While calibration focuses on representing the true correctness likelihood of ID data only, the ODIN score is designed to maximize the gap between ID and OOD data and may no longer be meaningful from a predictive confidence standpoint. Built on the insights, recent work [104, 111] proposed using an energy score for OOD detection, which enjoys theoretical interpretation from a likelihood perspective [293]. Test samples with lower energy are considered ID and vice versa. JointEnergy score [106] is then proposed to perform OOD detection for multi-label classification networks. Techniques such as layer-wise Mahalanobis distance [103] and Gram Matrix [105] are implemented for better hidden feature quality to perform density estimation.
Recently, one fundamental cause of the overconfidence issue on OOD data is revealed that using mismatched BatchNorm statistics—that are estimated on ID data yet blindly applied to the OOD data in testing—can trigger abnormally high unit activations and model output accordingly [107]. Therefore, ReAct [107] proposes truncating the high activations, which establishes strong post-hoc detection performance and further boosts the performance of existing scoring functions. Similarly, NMD [108] uses the activation means from BatchNorm layers for ID / OOD discrepancy. While ReAct considers activation space, [109] proposes a weight sparsification-based OOD detection framework termed DICE. DICE ranks weights based on a measure of contribution, and selectively uses the most salient weights to derive the output for OOD detection. By pruning away noisy signals, DICE provably reduces the output variance for OOD data, resulting in a sharper output distribution and stronger separability from ID data.
b. Confidence Enhancement Methods  Tailored for OOD detection problem, confidence can be developed via designing a confidence-estimating branch [112] or class [113], data augmentation [114], ensembling with leaving-out strategy [115], adversarial training [116, 117, 118, 119, 129], stronger data augmentation [120, 121, 122, 123], better uncertainty modeling [124, 294], and utilizing feature or statistics from the intermediate-layer features [111, 295]. Specially, to enhance the sensitivity to covariate shift, some methods focus on the hidden representations in the middle layers of neural networks. Generalized ODIN, or G-ODIN [69] extended ODIN [102] by using a specialized training objective termed DeConf-C and choose hyperparameters such as perturbation magnitude on ID data. Note that we do not categorize G-ODIN as post-hoc method as it requires model retraining. Recent work [125] shows that the overconfidence issue can be mitigated through Logit Normalization (LogitNorm), a simple fix to the common cross-entropy loss by enforcing a constant vector norm on the logits in training. Trained with LogitNorm, neural networks produce highly distinguishable confidence scores between in- and out-of-distribution data.
c. Outlier Exposure  Another branch of OOD detection methods makes use of a set of collected OOD samples, or “outlier”, during training to help models learn ID/OOD discrepancy. Starting from the concurrent baselines that encourage a flat/high-entropic prediction on given OOD samples [67, 59] and suppressing OOD feature magnitudes [59], a follow-up work, MCD [126] uses a network with two branches, between which entropy discrepancy is enlarged for OOD training data. Another straightforward approach with outlier exposure spares an extra abstention (or rejection class) and considers all the given OOD samples in this class [134, 129, 132]. A later work OECC [127] noticed that an extra regularization for confidence calibration introduces additional improvement for OE. To effectively utilize the given, usually massive, OOD samples, some works use outlier mining [129, 136] and adversarial resampling [128] approaches to obtain a compact yet representative set. Other works consider a more practical scenario where given OOD samples contain ID samples, therefore using pseudo-labeling [130] or ID filtering methods [131] to reduce the interference of ID data. In general, OOD detection with outlier exposure can reach a much better performance. However, as research shows that the performance can be largely affected by the correlations between given and real OOD samples [133]. To address the issue, recent work [135] proposes a novel framework that enables effectively exploiting unlabeled in-the-wild data for OOD detection. Unlabeled wild data is frequently available since it is produced essentially for free whenever deploying an existing classifier in a real-world system. This setting can be viewed as training OOD detectors in their natural habitats, which provide a much better match to the true test time distribution than data collected offline.
3.1.2 Label Space Redesign
One-hot encoding is commonly used to encode categorical information for classification. However, one-hot encoding ignores the inherent relationship among labels. For example, it is unreasonable to have a uniform distance between dog and cat vs. dog and car. To this end, several works attempt to use information in the label space for OOD detection. Some works arrange the large semantic space into a hierarchical taxonomy of known classes [137, 138]. Under the redesigned label architecture, top-down classification strategy [137] and group softmax training [138] are demonstrated effective. Another set of works uses word embeddings to automatically construct the label space. In [139], the sparse one-hot labels are replaced with several dense word embeddings from different NLP models, forming multiple regression heads for robust training. When testing, the label, which has the minimal distance to all the embedding vectors from different heads, will be considered as the prediction. If the minimal distance crosses above the threshold, the sample would be classified as “novel”. Recent works further take the image features from language-image pre-training models [296] to better detect novel classes, where the image encoding space also contains rich information from the language space [140, 141].
3.1.3 OOD Data Generation
The outlier exposure approaches impose a strong assumption on the availability of OOD training data, which can be infeasible in practice. When no OOD sample is available, some methods attempt to synthesize OOD samples to enable ID/OOD separability. Existing works leverage GANs to generate OOD training samples and force the model predictions to be uniform [142], generate boundary samples in the low-density region [143], or similarly, high-confidence OOD samples [144]. However, synthesizing images in the high-dimensional pixel space can be difficult to optimize. Recent work VOS [146] proposed synthesizing virtual outliers from the low-likelihood region in the feature space, which is more tractable given lower dimensionality. In object detection, [147] proposes synthesizing unknown objects from videos in the wild using spatial-temporal unknown distillation.
3.1.4 Gradient-based Methods
3.1.5 Bayesian Models
A Bayesian model is a statistical model that implements Bayes’ rule to infer all uncertainty within the model [297]. The most representative method is the Bayesian neural network [298], which draws samples from the posterior distribution of the model via MCMC [299], Laplace methods [300, 301] and variational inference [302], forming the epistemic uncertainty of the model prediction. However, their obvious shortcomings of inaccurate predictions [303] and high computational costs [304] prevent them from wide adoption in practice. Recent works attempt several less principled approximations including MC-dropout [149] and deep ensembles [305, 150, 306] for faster and better estimates of uncertainty. These methods are less competitive for OOD uncertainty estimation. Further exploration takes natural-gradient variational inference and enables practical and affordable modern deep learning training while preserving the benefits of Bayesian principles [151]. Dirichlet Prior Network (DPN) is also used for OOD detection with an uncertainty modeling of three different sources of uncertainty: model uncertainty, data uncertainty, and distributional uncertainty and form a line of works [152, 153, 154]. Recently, Bayesian hypothesis test is used to formulate OOD detection, with upweighting method and Hessian approximation for scalability [307].
3.1.6 Large-scale OOD Detection
3.2 Density-based Methods
Density-based methods in OOD detection explicitly model the in-distribution with some probabilistic models, and flag test data in low-density regions as OOD. Although OOD detection can be different from AD in that multiple classes exist in the in-distribution, density estimation methods used for AD in Section 5.1.2 can be directly adapted to OOD detection by unifying the ID data as a whole [157, 158, 159, 160, 161]. When the ID contains multiple classes, class-conditional Gaussian distribution can explicitly model the in-distribution so that the OOD samples can be identified based on their likelihoods [103]. Flow-based methods [162, 163, 164, 165, 166] can also be used for probabilistic modeling. While directly estimating the likelihood seems like a natural approach, some works [167, 168, 169] find that probabilistic models sometimes assign a higher likelihood for the OOD sample. Several works attempt to solve the problems using likelihood ratio [170].  [171] finds that the likelihood exhibits a strong bias towards the input complexity and proposes a likelihood ratio-based method to compensate the influence of input complexity. Recent methods turn to new scores such as likelihood regret [172] or an ensemble of multiple density models [168]. To directly model the density of semantic space, SEM score is used with a simple combination of density estimation in the low-level and high-level space [308]. Overall, generative models can be prohibitively challenging to train and optimize, and the performance can often lag behind the classification-based approaches (Section 3.1).
3.3 Distance-based Methods
The basic idea of distance-based methods is that the testing OOD samples should be relatively far away from the centroids or prototypes of in-distribution classes.  [103] uses the minimum Mahalanobis distance to all class centroids for detection. A subsequent work splits the images into foreground and background, and then calculates the Mahalanobis distance ratio between the two spaces [173]. In contrast to parametric approach, recent work [110] shows strong promise of non-parametric nearest-neighbor distance for OOD detection. Unlike Mahalanobis, the non-parametric approach does not impose any distributional assumption about the underlying feature space, hence providing stronger simplicity, flexibility and generality.
Some works use cosine similarity between test sample features and class features to determine OOD samples [174, 175]. The one-dimensional subspace spanned by the first singular vector of the training features is shown to be more suitable for cosine similarity-based detection [176]. Moreover, other works leverage distances with radial basis function kernel [177], Euclidean distance [178], and geodesic distance [309] between the input’s embedding and the class centroids. Apart from calculating the distance between samples and class centroids, feature norm that in the orthogonal complement space of the principal space is shown effective on OOD detection [310]. Distance-based OOD scores can generally benefit from enhanced representations. Recent work CIDER [179] introduces a new representation learning framework for OOD detection, which consists of a dispersion loss promoting large angular distances among different class prototypes, along with a compactness loss encouraging samples to be close to their class prototypes.
3.4 Reconstruction-based Methods
The core idea of reconstruction-based methods is that the encoder-decoder framework trained on the ID data usually yields different outcomes for ID and OOD samples. The difference in model performance can be utilized as an indicator for detecting anomalies. For example, reconstruction models that are only trained by ID data cannot well recover the OOD data [180], and therefore the OOD can be identified. While reconstruction-based models with pixel-level comparison seem not a popular solution in OOD detection for its expensive training cost, reconstructing with hidden features is shown as a promising alternative [181]. Rather than reconstructing the entire image, recent work MoodCat [182] masks a random portion of the input image and identifies OOD samples using the quality of the classification-based reconstruction results.
3.5 Discussion
The field of OOD detection has enjoyed rapid development since its emergence, with a large space of solutions. In the multi-class setting, the problem can be canonical to OSR (Section 4)—accurately classify test samples from ID within the class space $\mathcal{Y}$, and reject test samples with semantics outside the support of $\mathcal{Y}$. The difference often lies in the evaluation protocol. OSR splits a dataset into two halves: one set as ID and another set as OOD. In contrast, OOD allows a more general and flexible evaluation by considering test samples from different datasets or domains. Moreover, OOD detection encompasses a broader spectrum of learning tasks (e.g., multi-label classification [106], object detection [146, 147]) and solution space. Apart from the methodology development, theoretical understanding has also received attention in the community [293], providing provable guarantees and empirical analysis to understand how OOD detection performance changes with respect to data distributions.
4 Open Set Recognition: Methodology
4.1 Classification-based Methods
4.1.1 EVT-based Uncertainty Calibration
4.1.2 EVT-Free Confidence Enhancement
4.1.3 Unknown Class Generation
4.2 Distance-based Methods
Distance-based methods for OSR require the prototypes to be class-conditional, which allows maintaining the ID classification performance. Category-based clustering and prototyping are performed based on the visual features extracted from the classifiers. OOD samples can be detected by computing the distance w.r.t. clusters [197, 198]. Some methods also leveraged contrastive learning to learn more compact clusters for known classes [199, 200], which enlarge the distance between ID and OOD. CROSR [201] enhances the features by concatenating visual embeddings from both the classifier and reconstruction model for distance computation in the extended feature space. Besides using features from classifiers, GMVAE [202] extracts features using a reconstruction VAE, and models the embeddings of the training set as a Gaussian mixture with multiple centroids for the following distance-based operations. Classifiers using nearest neighbors are also adapted for OSR problem [203]. By storing the training samples, the nearest neighbor distance ratio is used for identifying unknown samples in testing.
4.3 Reconstruction-based Methods
4.3.1 Sparse Representation Methods
4.3.2 Reconstruction-Error Methods
4.4 Discussion
Due to the restriction on using only ID data for training, OSR methods do not implement background classes, or outlier exposure (c.f. Section 3.1.1). A recent research shows that a good classifier [314] in the close-set is comparable to the state-of-the-art OSR methods, which urges the community to develop more general approaches without great efforts on carving for specific datasets.
5 Anomaly Detection & Novelty Detection: Methodology